# fev-bench: A Realistic Benchmark for Time Series Forecasting

### v1

@@ 6.1 Existing benchmarks
Early works on deep learning approaches for time series forecasting (Lai et al., 2018; Salinas et al., 2020; Rangapuram et al., 2018) did not rely on standardized benchmarks. Instead, they typically evaluated on 4–6 datasets from different domains, which often varied across studies. The M4 (Makridakis et al., 2020) and M5 (Makridakis et al., 2022) competition tasks gained popularity for evaluating forecasting models. The LTSF benchmark (Zhou et al., 2021; Zeng et al., 2023) introduced new datasets with an emphasis on long-horizon forecasting. However, these benchmarks were limited in scope, focusing on narrow domains without the comprehensiveness needed to assess general forecasting models. In addition, the LTSF benchmark has been criticized for its over-representation of similar datasets and impractical evaluation tasks (Hewamalage et al., 2023).
A subsequent wave of benchmarks, including the Monash forecasting repository (Godahewa et al., 2021), BasicTS+ (Shao et al., 2024), TFB (Qiu et al., 2024), ProbTS (Zhang et al., 2024), and Chronos Benchmarks (Ansari et al., 2024), broadened evaluation by covering datasets from diverse domains and supporting both univariate and multivariate forecasting. More recently, GIFT-Eval (Aksu et al., 2024) greatly expanded the range of forecasting tasks with varied domains and frequencies, quickly becoming the standard benchmark for pretrained time series models. Domain-specific benchmarks, such as BOOM (Cohen et al., 2025), have also been introduced. Yet, none of these benchmarks include forecasting tasks with covariates, despite their immense practical relevance.
In contrast, fev-bench provides broader dataset and domain coverage, including a substantial proportion of tasks with covariates, along with domain-appropriate forecasting setups. Complementing this, the fev library offers standardized infrastructure for evaluation, incorporating principled aggregation strategies and confidence intervals, leading to more robust conclusions.
@@ 6.2 Aggregation strategies
Several aggregation methods have been proposed in the literature, each with specific trade-offs for benchmarking.
Average rank has been widely used in the benchmarking literature (Aksu et al., 2024; Ansari et al., 2024). As we show in Section C.1, average rank is mathematically equivalent to the average win rate $W_{j}$. The two induce exactly the same ordering of models, differing only by an affine transformation. Since ranks scale with the number of models and lack a natural extension to pairwise comparisons, we prefer reporting win rates, which are bounded between 0 and $1$ and extend directly to pairwise evaluations.
Model Avg. win rate (%) Skill score (%) Median runtime (s) Leakage (%) # failures
TiRex 86.7 42.6 1.4 1 0
TimesFM-2.5 82.1 42.3 117.6 8 0
Toto-1.0 73.8 40.7 90.7 8 0
Moirai-2.0 68.8 39.3 2.5 28 0
Chronos-Bolt 68.8 38.9 1.0 0 0
TabPFN-TS 66.9 39.6 305.5 0 2
Sundial-Base 49.2 33.4 35.6 1 0
Stat. Ensemble 48.7 20.2 690.6 0 11
Seasonal Naive 21.7 0.0 2.3 0 0
@@
Model Avg. win rate (%) Skill score (%) Median runtime (s) Leakage (%) # failures
TiRex 80.5 30.0 1.4 1 0
TimesFM-2.5 79.9 30.3 117.6 8 0
Toto-1.0 69.9 28.2 90.7 8 0
Moirai-2.0 65.2 27.3 2.5 28 0
Chronos-Bolt 64.8 26.5 1.0 0 0
TabPFN-TS 62.0 27.6 305.5 0 2
Sundial-Base 56.7 24.7 35.6 1 0
Stat. Ensemble 51.0 15.7 690.6 0 11
Seasonal Naive 22.3 0.0 2.3 0 0
@@
Bradley–Terry (Elo) scores (Bradley & Terry, 1952) have been applied in diverse domains, from large language models (Chiang et al., 2024) to tabular benchmarks (Erickson et al., 2025). As we show in Section C.2, when all models are compared on all tasks, Bradley–Terry scores induce the same model ranking as average win rates. In other words, average rank, average win rate, and Bradley–Terry scores are equivalent in terms of model ordering given our setup in Section 4. We choose to report win rates for simplicity.
Nemenyi post-hoc test with critical difference diagrams (Demšar, 2006) is another rank-based method, since it relies on average ranks as input. It controls family-wise error rates but is often overly conservative, frequently failing to detect meaningful differences (Garcia & Herrera, 2008). Moreover, it yields only binary significance decisions without quantifying effect sizes. Confidence intervals on win rates provide a more informative alternative: they preserve the intuitive interpretation of win rates and display the magnitude and uncertainty of performance differences directly.
Geometric mean relative error (GMRE) (Ansari et al., 2024; Aksu et al., 2024) yields rankings identical to skill scores since $\operatorname{GMRE}_{j}=1-S_{j}$. We adopt this established approach with two modifications: extreme values are clipped to limit the contribution of outliers, and we report its complement (the skill score) to maintain a consistent “higher-is-better” interpretation across all aggregation methods.
@@ 7 Results
Figure 1: Pairwise win rates (a) and skill scores (b) of the top-3 models against other models under the SQL metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better. The confidence intervals show heavy overlap between TiRex and TimesFM-2.5, suggesting no clear winner between the two, while both outperform the remaining models. Full pairwise results are available in Appendix D. Best viewed on screen.
@@ 7.1 Setup
We evaluate seven pretrained models: TimesFM-2.5 (Das et al., 2024), TiRex (Auer et al., 2025), Chronos-Bolt (Base) (Ansari et al., 2024), Toto-1.0 (Cohen et al., 2025), Moirai-2.0 (Woo et al., 2024), TabPFN-TS (Hoo et al., 2025), and Sundial (Liu et al., 2025). Among these models only Toto-1.0 natively supports multivariate forecasting ($D>1$). For the remaining tasks, we first convert each multivariate time series in the original dataset into $D$ separate univariate series, one for each dimension. TabPFN-TS is the only model that supports known covariates, the remaining models ignore all covariates.
We include statistical baselines representing different modeling approaches: AutoETS, AutoARIMA, and AutoTheta (Garza et al., 2022), plus SCUM ensemble that combines the aforementioned univariate models with AutoCES (Petropoulos & Svetunkov, 2020). We also evaluate simple baselines Seasonal Naive, Naive, and Drift (Hyndman & Athanasopoulos, 2018).
@@
Evaluation metrics. We follow the evaluation protocol described in Sections 3 and 4. Each task is evaluated using SQL for probabilistic forecasting and MASE for point forecasting, with results aggregated across all 100 tasks using average win rates and skill scores for both marginal and pairwise model comparisons. In addition, we report the following metrics to provide broader context for model performance.
Data leakage. The purpose of fev-bench is to assess the zero-shot capabilities of pretrained forecasting models. Two kinds of leakage can undermine this goal: (i) if a model has been trained on the benchmark training split, the evaluation is no longer zero-shot; and (ii) if the model has seen the test split, this constitutes direct test leakage. To prevent both, model contributors must indicate, for each model–task pair, whether any part of the dataset at the same frequency was used during training. Resampled variants at different frequencies are not considered leakage. For tasks where overlap is reported, we discard the submitted results and impute performance with the 100% zero-shot model that had the highest average win rate at the time of the benchmark release, namely Chronos-Bolt (Base). This eliminates leakage for the affected tasks without heavily penalizing the submitted model.
@@
Runtime. In addition to accuracy indicators, we report the median end-to-end runtime (training plus prediction across all evaluation windows) for each model. While this metric has limitations due to varying hardware, batch sizes, and implementation details, it still offers useful insights into computational efficiency. In practice, it helps distinguish between different modeling paradigms (e.g., sample-based autoregressive models versus direct multi-step forecasters) and between zero-shot and fine-tuned approaches, while also providing an incentive for contributors to develop and submit more efficient implementations.
@@ 7.2 Results
Tables 4 & 5 summarize the marginal performance of the top-performing models on fev-bench. Complete marginal and pairwise results for all models are provided in Appendix D. Live results for the leaderboard are available on Hugging Face.333huggingface.co/spaces/autogluon/fev-leaderboard
The overall ranking aligns with findings from other benchmarks (Aksu et al., 2024; Ansari et al., 2024). TiRex and TimesFM-2.5 emerge as the top two models, leading the benchmark in both point and probabilistic forecast accuracy. Older pretrained models such as Toto-1.0, Chronos-Bolt and Sundial rank lower. All pretrained models are substantially more accurate and faster than the best statistical approach (SCUM Ensemble).
To assess whether the observed differences among the leading models reflect genuine improvements rather than evaluation noise, we examine pairwise comparisons under the SQL metric in Figures 1a and 1b. The confidence intervals show no statistically significant gap between TiRex and TimesFM-2.5, despite the difference in their marginal win rates and skill scores in Table 4. This indicates that under different benchmark compositions or task weightings, either of the two could emerge as the top performer. In contrast, both TiRex and TimesFM-2.5 demonstrate clear advantages over all other models, with confidence intervals that separate them from the rest of the field.
Figure 2: Average skill scores on the 42 fev-bench tasks with dynamic covariates (based on SQL). TabPFN-TS, the only model that uses known covariates, outperforms all others, indicating that pretrained models miss valuable predictive signal from covariates.
Figure 3: Average skill scores relative to the baseline on 35 multivariate tasks of fev-bench (based on SQL). Toto-1.0, the only multivariate model, outperforms others despite ranking third overall in Table 4, indicating room for improvement on multivariate forecasting.
@@ 7.3 Directions for improvement
Forecasting with covariates. On the 42 tasks in fev-bench with dynamic covariates, we compared the same set of pretrained models as in the main evaluation. As shown in Figure 2, TabPFN-TS achieves the highest skill score on covariate tasks (40.0%), clearly ahead of the second-best model TiRex (37.8%). This marks a notable change compared to the overall results (Table 4), where TabPFN-TS ranked fourth by skill score. This demonstrates that current pretrained models leave substantial performance untapped by ignoring covariates.
Multivariate forecasting. Restricting attention to the 35 multivariate tasks in fev-bench, Toto-1.0 outperforms TiRex, reversing the overall ranking (Figure 3). Toto-1.0 is the only model that natively supports multivariate forecasting; all others predict each dimension independently. This advantage underscores the need for pretrained models that handle multivariate series directly, though doing so requires advances in both architecture design and access to multivariate training data.
@@ Appendix A Tasks
The precise task definitions in YAML format are available under https://github.com/autogluon/fev. The datasets used for evaluation are hosted on Hugging Face https://huggingface.co/datasets/autogluon/fev_datasets.
@@ Appendix B Models
We evaluate seven pretrained models on fev-bench, whose key properties are summarized in Table 12. Most are decoder-only transformers, except TiRex, a decoder-only xLSTM model, and Chronos-Bolt (Base), an encoder-decoder transformer. All models except TabPFN-TS process non-overlapping patches of time series rather than individual observations. Toto, TabPFN-TS, and Sundial produce sample forecasts, while the others generate quantiles on a fixed grid. Toto is the only model that natively supports multivariate targets. TabPFN-TS is the only model that accepts known covariates, though we did not provide past covariates since they consistently degraded its accuracy.
@@ Appendix C Extended discussion of aggregation methods
This appendix clarifies how average win rates $W_{j}$ (Equation 6) relate to other aggregation methods commonly used in benchmarking, such as average ranks (Aksu et al., 2024; Ansari et al., 2024) and Bradley–Terry (“Elo”) scores (Erickson et al., 2025). We show that all three induce the same ordering of models.
@@ C.2 Average win rate and Bradley–Terry (Elo) scores result in the same ranking
Suppose all $M$ models are compared on the same $R$ tasks, with pairwise win rates $W_{jk}$ (Equation 8) and average win rates $W_{j}=\tfrac{1}{M-1}\sum_{k\neq j}W_{jk}$ (Equation 6). At the BT MLE with scale $\lambda>0$,
@@ Appendix D Extended results
Tabular results for each dataset-task combination, updated marginal and pairwise results in an interactive foramt are available under https://huggingface.co/spaces/autogluon/fev-leaderboard.
@@ D.1 Marginal performance
Model Avg. win rate (%) Skill score (%) Median runtime (s) Leakage (%) # failures
TiRex 86.7 42.6 1.4 1 0
TimesFM-2.5 82.1 42.3 117.6 8 0
Toto-1.0 73.8 40.7 90.7 8 0
Moirai-2.0 68.8 39.3 2.5 28 0
Chronos-Bolt 68.8 38.9 1.0 0 0
TabPFN-TS 66.9 39.6 305.5 0 2
Sundial-Base 49.2 33.4 35.6 1 0
Stat. Ensemble 48.7 20.2 690.6 0 11
AutoARIMA 43.5 20.6 186.8 0 10
AutoETS 35.8 -26.8 17.0 0 3
AutoTheta 29.2 5.5 9.3 0 0
Seasonal Naive 21.7 0.0 2.3 0 0
Naive 14.9 -45.4 2.2 0 0
Drift 9.9 -45.8 2.2 0 0
@@
Model Avg. win rate (%) Skill score (%) Median runtime (s) Leakage (%) # failures
TiRex 80.5 30.0 1.4 1 0
TimesFM-2.5 79.9 30.3 117.6 8 0
Toto-1.0 69.9 28.2 90.7 8 0
Moirai-2.0 65.2 27.3 2.5 28 0
Chronos-Bolt 64.8 26.5 1.0 0 0
TabPFN-TS 62.0 27.6 305.5 0 2
Sundial-Base 56.7 24.7 35.6 1 0
Stat. Ensemble 51.0 15.7 690.6 0 11
AutoARIMA 39.0 11.2 186.8 0 10
AutoTheta 37.1 11.0 9.3 0 0
AutoETS 34.9 2.3 17.0 0 3
Seasonal Naive 22.3 0.0 2.3 0 0
Naive 20.6 -16.7 2.2 0 0
Drift 16.0 -18.1 2.2 0 0

### v2

@@ 6.1 Existing benchmarks
Early deep learning approaches for time series forecasting (Lai et al., 2018; Salinas et al., 2020; Rangapuram et al., 2018) did not rely on standardized benchmarks, instead evaluating on a small number of datasets that varied across studies. The M4 (Makridakis et al., 2020) and M5 (Makridakis et al., 2022) competition tasks later gained popularity. The LTSF benchmark (Zhou et al., 2021; Zeng et al., 2023) introduced new datasets with an emphasis on long-horizon forecasting, but remained limited in scope, with a narrow domain focus. It has also been criticized for over-representing similar datasets and using impractical evaluation tasks (Hewamalage et al., 2023).
Subsequent benchmarks, including the Monash repository (Godahewa et al., 2021), BasicTS+ (Shao et al., 2024), TFB (Qiu et al., 2024), ProbTS (Zhang et al., 2024), and Chronos Benchmarks (Ansari et al., 2024), broadened evaluation by covering more domains and supporting both univariate and multivariate forecasting. More recently, GIFT-Eval (Aksu et al., 2024) substantially expanded task diversity across domains and frequencies, becoming a standard benchmark for pretrained time series models. Domain-specific benchmarks such as BOOM (Cohen et al., 2025) have also been introduced. However, none of these benchmarks include forecasting tasks with covariates, despite their practical importance.
In contrast, fev-bench offers broader dataset and domain coverage, including a substantial number of tasks with covariates and domain-appropriate forecasting setups. Complementing this, the fev library provides standardized evaluation infrastructure with principled aggregation and confidence intervals, enabling more robust comparisons.
@@ 6.2 Aggregation strategies
Several aggregation methods have been proposed in the literature, each with distinct trade-offs for benchmarking.
Average rank is widely used in forecasting benchmarks (Aksu et al., 2024; Ansari et al., 2024). As shown in Sec. C.1, average rank is mathematically equivalent to the average win rate $W_{j}$, inducing the same model ordering. However, ranks scale with the number of models and do not naturally extend to pairwise comparisons. We therefore prefer win rates, which are bounded between 0 and $1$ and extend directly to pairwise evaluation.
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 91.4 47.3 0.8 0 0
TiRex 82.7 42.6 0.2 1 0
TimesFM-2.5 77.6 42.2 1.9 10 0
Toto-1.0 70.2 40.7 22.1 8 0
Moirai-2.0 64.4 39.3 0.3 28 0
Chronos-Bolt 64.4 38.9 0.2 0 0
TabPFN-TS 63.0 39.6 88.9 0 2
Sundial-Base 46.0 33.4 8.0 1 0
Stat. Ensemble 45.5 20.2 148.6 0 11
Seasonal Naive 20.2 0.0 0.5 0 0
@@
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 88.1 35.5 0.8 0 0
TiRex 76.9 30.0 0.2 1 0
TimesFM-2.5 75.1 30.2 1.9 10 0
Toto-1.0 66.9 28.2 22.1 8 0
Moirai-2.0 61.2 27.3 0.3 28 0
Chronos-Bolt 60.8 26.5 0.2 0 0
TabPFN-TS 58.7 27.6 88.9 0 2
Sundial-Base 53.4 24.7 8.0 1 0
Stat. Ensemble 48.5 15.7 148.6 0 11
Seasonal Naive 21.0 0.0 0.5 0 0
@@
Bradley–Terry (Elo) scores (Bradley & Terry, 1952) have been applied across domains, from large language models (Chiang et al., 2024) to tabular benchmarks (Erickson et al., 2025). As shown in Sec. C.2, when all models are compared on all tasks, Bradley–Terry scores induce the same model ranking as average win rates. Thus, average rank, average win rate, and Bradley–Terry scores are equivalent in terms of model ordering under our setup (Sec. 4). We report win rates for simplicity.
Nemenyi post-hoc test with critical difference diagrams (Demšar, 2006) is another rank-based method, as it relies on average ranks. While it controls family-wise error rates, it is often overly conservative and frequently fails to detect meaningful differences (Garcia & Herrera, 2008). Moreover, it provides only binary significance decisions without effect sizes. In contrast, confidence intervals on win rates preserve interpretability while directly conveying the magnitude and uncertainty of performance differences.
Geometric mean relative error (GMRE) (Ansari et al., 2024; Aksu et al., 2024) yields rankings identical to skill scores since $\operatorname{GMRE}_{j}=1-S_{j}$. We adopt this approach with two modifications: clipping extreme values to limit outlier influence, and reporting its complement (the skill score) to maintain a consistent “higher-is-better” interpretation.
@@ 7 Results
Figure 1: Pairwise win rates (a) and skill scores (b) of the top-4 models against other models under the SQL metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better. Full pairwise results are available in Appendix D. Best viewed on screen.
@@ 7.1 Setup
We evaluate eight pretrained models: Chronos-2 (Ansari et al., 2025), TimesFM-2.5 (Das et al., 2024), TiRex (Auer et al., 2025), Chronos-Bolt (Base) (Ansari et al., 2024), Toto-1.0 (Cohen et al., 2025), Moirai-2.0 (Woo et al., 2024), TabPFN-TS (Hoo et al., 2025), and Sundial (Liu et al., 2025). Chronos-2 and Toto-1.0 natively support multivariate forecasting ($D>1$); for the remaining models, we convert each multivariate time series into $D$ separate univariate series. Chronos-2 supports both past and future covariates, TabPFN-TS supports known covariates, and the remaining models ignore all covariates.
We include popular statistical baselines: AutoETS, AutoARIMA, and AutoTheta (Garza et al., 2022), as well as the SCUM ensemble combining these univariate models with AutoCES (Petropoulos & Svetunkov, 2020). We also evaluate simple baselines: Seasonal Naive, Naive, and Drift (Hyndman & Athanasopoulos, 2018).
@@
Evaluation metrics. We follow the evaluation protocol described in Sections 3 and 4. Each task is evaluated using SQL for probabilistic forecasting and MASE for point forecasting, with results aggregated across all 100 tasks using average win rates and skill scores for both marginal and pairwise model comparisons. We also report the following metrics to provide broader context for model performance.
Data leakage. The goal of fev-bench is to assess the zero-shot capabilities of pretrained forecasting models. Two forms of leakage undermine this goal: (i) training on the benchmark training split, which breaks the zero-shot setting, and (ii) exposure to the test split, which constitutes direct test leakage. To prevent both, contributors must indicate, for each model–task pair, whether any part of the dataset at the same frequency was used during training. Resampled variants at different frequencies are not considered leakage. For tasks with reported overlap, we discard the submitted results and impute performance using the strongest fully zero-shot univariate model at release time, Chronos-Bolt (Base). This removes leakage for affected tasks without heavily penalizing the submitted model.
@@
Runtime. In addition to accuracy indicators, we report the median end-to-end runtime (training plus inference time across all evaluation windows) for each model. Runtimes are normalized to 100 time series per task by scaling each task’s runtime by $100/(N\times W)$, where $N$ is the number of items and $W$ the number of evaluation windows. While this metric is affected by hardware, batch sizes, and implementation details, it still provides useful insights into computational efficiency. In practice, it helps distinguish between modeling paradigms (e.g., sample-based autoregressive models versus direct multi-step forecasters) and between zero-shot and fine-tuned approaches, while incentivizing more efficient implementations.
@@ 7.2 Results
Tables 4 & 5 summarize the marginal performance of the top-performing models on fev-bench. Complete results for all models are provided in Appendix D. Chronos-2 achieves the highest average win rate and skill score values, followed by TiRex and TimesFM-2.5. Older pretrained models such as Chronos-Bolt and Sundial rank lower. All pretrained models are substantially more accurate and faster than the best statistical approach (SCUM Ensemble).
To assess whether the observed differences among the leading models reflect genuine improvements rather than evaluation noise, we examine pairwise comparisons under the SQL metric in Figures 1a and 1b. The confidence intervals indicate a statistically significant gap between Chronos-2 and the remaining models. In contrast, TiRex and TimesFM-2.5 show heavy overlap, suggesting no clear winner between the two; under different benchmark compositions or task weightings, either could emerge as the second-best performer.
Figure 2: Average skill scores on the 42 fev-bench tasks with dynamic covariates (based on SQL). Chronos-2 and TabPFN-TS, the only models that use covariates, outperform others on these tasks.
@@ 7.3 Directions for improvement
Figure 3: Average skill scores relative to the baseline on 35 multivariate tasks of fev-bench (based on SQL). Chronos-2 and Toto-1.0, the only models with native multivariate support, perform well on these tasks.
Forecasting with covariates. On the 42 tasks in fev-bench with dynamic covariates, we compared the same set of pretrained models as in the main evaluation. As shown in Figure 2, models that support covariates (Chronos-2 and TabPFN-TS) achieve higher skill scores on these tasks compared to models that ignore covariates. This demonstrates that covariate support provides meaningful performance gains when such information is available.
Multivariate forecasting. Restricting attention to the 35 multivariate tasks in fev-bench, models with native multivariate support (Chronos-2 and Toto-1.0) perform well (Figure 3). Other models predict each dimension independently. This highlights the benefit of modeling cross-variate dependencies when forecasting multivariate time series, and underscores the need for pretrained models that handle multivariate series directly, though doing so requires advances in both architecture design and access to multivariate training data.
@@ Appendix A Tasks
The precise task definitions in YAML format are available under https://github.com/autogluon/fev. The datasets used for evaluation are available under https://huggingface.co/datasets/autogluon/fev_datasets.
@@ Appendix B Models
We evaluate seven pretrained models on fev-bench, whose key properties are summarized in Table 12. Most are decoder-only transformers, except TiRex, a decoder-only xLSTM model, Chronos-Bolt (Base), an encoder-decoder transformer, and Chronos-2, an encoder-only model. All models except TabPFN-TS process non-overlapping patches of time series rather than individual observations. Toto, TabPFN-TS, and Sundial produce sample forecasts, while the others generate quantiles on a fixed grid. Toto and Chronos-2 are the only models that natively supports multivariate targets. TabPFN-TS supports known covariates, though we did not provide past covariates since they consistently degraded its accuracy. Chronos-2 supports both past and known covariates.
@@
Chronos-2 amazon/chronos-2 100 2.2.2 8192 {cross_learning: True}
@@ Appendix C Extended discussion of aggregation methods
This appendix clarifies how average win rates $W_{j}$ (Eq. 4) relate to other aggregation methods commonly used in benchmarking, such as average ranks (Aksu et al., 2024; Ansari et al., 2024) and Bradley–Terry (“Elo”) scores (Erickson et al., 2025). We show that all three induce the same ordering of models.
@@ C.2 Average win rate and Bradley–Terry (Elo) scores result in the same ranking
Suppose all $M$ models are compared on the same $R$ tasks, with pairwise win rates $W_{jk}$ (Eq. 6) and average win rates $W_{j}=\tfrac{1}{M-1}\sum_{k\neq j}W_{jk}$ (Eq. 4). At the BT MLE with scale $\lambda>0$,
@@ Appendix D Extended results
Tabular results for each dataset-task combination, updated marginal and pairwise results in an interactive format are available under https://huggingface.co/spaces/autogluon/fev-bench.
@@ D.1 Marginal performance
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 91.4 47.3 0.8 0 0
TiRex 82.7 42.6 0.2 1 0
TimesFM-2.5 77.6 42.2 1.9 10 0
Toto-1.0 70.2 40.7 22.1 8 0
Moirai-2.0 64.4 39.3 0.3 28 0
Chronos-Bolt 64.4 38.9 0.2 0 0
TabPFN-TS 63.0 39.6 88.9 0 2
Sundial-Base 46.0 33.4 8.0 1 0
Stat. Ensemble 45.5 20.2 148.6 0 11
AutoARIMA 40.7 20.6 19.5 0 10
AutoETS 33.7 -26.8 3.5 0 3
AutoTheta 27.1 5.5 3.3 0 0
Seasonal Naive 20.2 0.0 0.5 0 0
Naive 13.8 -45.4 0.5 0 0
Drift 9.3 -45.8 0.5 0 0
@@
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 88.1 35.5 0.8 0 0
TiRex 76.9 30.0 0.2 1 0
TimesFM-2.5 75.1 30.2 1.9 10 0
Toto-1.0 66.9 28.2 22.1 8 0
Moirai-2.0 61.2 27.3 0.3 28 0
Chronos-Bolt 60.8 26.5 0.2 0 0
TabPFN-TS 58.7 27.6 88.9 0 2
Sundial-Base 53.4 24.7 8.0 1 0
Stat. Ensemble 48.5 15.7 148.6 0 11
AutoARIMA 36.7 11.2 19.5 0 10
AutoTheta 34.9 11.0 3.3 0 0
AutoETS 33.2 2.3 3.5 0 3
Seasonal Naive 21.0 0.0 0.5 0 0
Naive 19.3 -16.7 0.5 0 0
Drift 15.4 -18.1 0.5 0 0